The Magellanic Clouds Newsletter

An electronic exchange on Magellanic Clouds research

Edited by Eva K. Grebel and You-Hua Chu

                                    mcnews@astro.uiuc.edu
                    http://www.astro.uiuc.edu/projects/mcnews/MCNews.html
                            http://www.astro.uni-bonn.de/~mcnews/

No. 38 November 15, 1999

Contents

News  1
Abstracts of refereed papers 2
Conference contributions  9
Thesis abstract  11
Job opportunity  11

News

Proceedings of IAU Symposium 190

We are pleased to announce that we have finished editing the proceedings of the IAU Symposium 190, “New Views of the Magellanic Clouds”. The hardcopy manuscripts have been delivered to the publisher, the Astronomical Society of the Pacific, and corrections have been made. We were told that this volume would be published in 1999. The table of contents can be found at

http://cadcwww.hia.nrc.ca/iau190/.

We apologize for the delay in producing these proceedings. While some important papers indeed arrived late, other responsibilities prevented the prompt attention the proceedings deserved.

To ensure that the proceedings are mailed to you correctly, please send e-mail to David Bohlender at david.bohlender@HIA.NRC.CA to update your postal address, if you have moved since the meeting was held in July 1998.

Thank you.

You-Hua Chu, Nicholas Suntzeff, Jim Hesser, and David Bohlender

Abstracts of Refereed Papers

Radiation hydrodynamics of SN 1987A: I. Global analysis of the light curve for the first 4 months

S. Blinnikov^{1}, P. Lundqvist^{2}, O. Bartunov^{1}, K. Nomoto^{3}, K. Iwamoto^{3}

1 Institute for Theoretical and Experimental Physics, 117259, Moscow, Russia 2 Stockholm Observatory, SE-133 36 Saltsjöbaden, Sweden 3 Department of Astronomy and the Research Center for the Early Universe, University of Tokyo, Tokyo 113-0033, Japan

The optical/UV light curves of SN 1987A are analyzed with the multi-energy group radiation hydrodynamics code stella. The calculated monochromatic and bolometric light curves are compared with observations shortly after shock breakout, during the early plateau, through the broad second maximum, and during the earliest phase of the radioactive tail. We have concentrated on a progenitor model calculated by Nomoto & Hashimoto and Saio, Nomoto, & Kato, which assumes that 14 M_{⊙} of the stellar mass is ejected. Using this model, we have updated constraints on the explosion energy and the extent of mixing in the ejecta. In particular, we determine the most likely range of E/M (explosion energy over ejecta mass) and R_{0} (radius of the progenitor). In general, our best models have energies in the range E = (1.1 ± 0.3) × 10^{51} ergs, and the agreement is better than in earlier, flux-limited diffusion calculations for the same explosion energy. Our modeled B and V fluxes compare well with observations, while the flux in U undershoots after ∼ 10 days by a factor of a few, presumably due to NLTE and line transfer effects. We also compare our results with IUE observations, and a very good quantitative agreement is found for the first days, and for one IUE band (2500 − 3000 Å) as long as for 3 months. We point out that the V flux estimated by McNaught & Zoltowski should probably be revised to a lower value.

Accepted by: The Astrophysical Journal (Main Journal) For preprints, contact peter@astro.su.se

Also available from the URL http://babbage.sissa.it/abs/astro-ph/9911205

An Unprecedented Change in the Spectrum of S Doradus: As Cool As It Gets

Philip Massey^{1}

1 Kitt Peak National Observatory, National Optical Astronomy Observatories, USA

Recent observations of S Doradus, the archetype of Luminous Blue Variables (LBVs), reveal that the star’s optical spectrum now resembles an F-type supergiant, with a rich complex of absorption lines. Despite nearly 50 years of spectroscopic monitoring, such a spectrum has never before been seen for S Dor despite numerous occasions when the star was equally bright. However, such F-type spectra have been seen in other LBVs, including Var B in M33 during a recent outburst, and in Eta Carina during an outburst in 1893. The singly ionized metal lines arise in a layer moving away from the star (towards us) at 50 km s^{−1}, consistent with the lines forming in a “pseudo photosphere” originating in the stellar wind. The temperature suggested by the F-type spectrum is as cool as an LBV can get. Our CCD UBV photometry shows that the star has brightened only 0.3 mag in V since 1996, at which time it showed an emission-line spectrum characteristic of LBVs at minimum. The de-reddened B −V color is consistent with an effective temperature suggested by the spectrum (7500^{∘}K), but there is a pronounced UV excess as evidenced by U − B. Why the behavior of S Dor has suddenly changed remains a mystery.

Accepted by: Publications of the Astronomical Society of the Pacific For preprints, contact massey@noao.edu

Also available from the URL ftp://gemini.tuc.noao.edu/pub/massey/Sdor.ps
or by anonymous ftp at gemini.tuc.noao.edu/pub/massey/Sdor.ps

ASCA Discovery of a Be X-Ray Pulsar in the SMC: AX J0051−733

Kensuke Imanishi^{1}, Jun yokogawa^{1}, Masahiro Tsujimoto^{1}, and Katsuji Koyama^{1,2}

1 Department of Physics, Graduate School of Science, Kyoto University, Sakyo-ku, Kyoto, 606-8502, Japan 2 CREST, Japan Science and Technology Corporation (JST), 4-1-8 Honmachi, Kawaguchi, Saitama 332-0012, Japan

ASCA observed the central region of the Small Magellanic Cloud, and found a hard X-ray source, AX J0051−733, at the position of the ROSAT source RX J0050.8−7316, which has an optical counterpart of a Be star. Coherent X-ray pulsations of 323.1 ± 0.3 s were discovered from AX J0051−733. The pulse profile shows several sub-peaks in the soft (0.7–2.0 keV) X-ray band, but becomes nearly sinusoidal in the harder (2.0–7.0 keV) X-ray band. The X-ray spectrum was found to be hard, and is well fitted by a power-law model with a photon index of 1.0 ± 0.4. The long-term flux history was examined with the archival data of Einstein observatory and ROSAT; a flux variability with a factor ≥ 10 was found.

Accepted by: Publications of the Astronomical Society of Japan

For preprints, contact kensuke@cr.scphys.kyoto-u.ac.jp
Also available from the URL http://www-cr.scphys.kyoto-u.ac.jp/MC/
or http://xxx.lanl.gov/abs/astro-ph/9910463

Discovery of a New X-Ray Pulsar, AX J0049−729, in the Small Magellanic Cloud with ASCA

Jun Yokogawa^{1}, Kensuke Imanishi^{1}, Masahiro Tsujimoto^{1}, Makoto Kohno^{1} and Katsuji Koyama^{1,2}

1 Department of Physics, Graduate School of Science, Kyoto University, Sakyo-ku, Kyoto, 606-8502, Japan 2 CREST, Japan Science and Technology Corporation (JST), 4-1-8 Honmachi, Kawaguchi, Saitama 332-0012, Japan

We report on the results of two ASCA observations on an X-ray source, AX J0049−729 (= RX J0049.1−7250), in the Small Magellanic Cloud. A sinusoidal pulsation with a barycentric period of 74.676 ± 0.022 s was found. The pulse fraction was ∼ 70%, independent of the X-ray energy. The spectrum was hard with an absorbed power-law having a photon index of ∼ 1. The X-ray flux was 1.1 × 10^{−11} erg s^{−1} cm^{−2}, and then declined to 1.6 × 10^{−12} erg s^{−1} cm^{−2} from the first to the second observations, separated by about one month. Together with the previous results of the Einstein and ROSAT satellites, we found that AX J0049−729 has been in quiescent states, but occasionally showing flare activities. The flux variability, long-pulse period, and hard spectrum support the idea that AX J0049−729 is an X-ray binary pulsar, most likely a Be star binary.

Accepted by: Publications of the Astronomical Society of Japan (vol. 51, 547)

For preprints, contact jun@cr.scphys.kyoto-u.ac.jp
Also available from the URL http://www-cr.scphys.kyoto-u.ac.jp/member/jun/job/

Discovery of an X-Ray Pulsar in the SMC: AX J0058−7203

Masahiro Tsujimoto^{1}, Kensuke Imanishi^{1}, Jun Yokogawa^{1} and Katsuji Koyama^{1,2}

1 Department of Physics, Graduate School of Science, Kyoto University, Sakyo-ku, Kyoto, 606-8502, Japan 2 CREST, Japan Science and Technology Corporation (JST), 4-1-8 Honmachi, Kawaguchi, Saitama 332-0012, Japan

We report on the discovery and analysis of an X-ray pulsar, AX J0058−7203, in the Small Magellanic Cloud. This pulsar exhibits coherent pulsations at P = 280.3 s with a double-peak structure. The X-ray spectrum is well fitted with a simple power-law model of photon index Γ ∼ 0.7. No significant change of the pulsation period over the observation was found. A comparison with ROSAT observations in the same field reveals that AX J0058−7203 is highly variable, and is most likely a Be star binary pulsar.

Accepted by: Publications of the Astronomical Society of Japan

For preprints, contact tsujimot@cr.scphys.kyoto-u.ac.jp
Also available from the URL http://www-cr.scphys.kyoto-u.ac.jp/MC/
or http://xxx.lanl.gov/abs/astro-ph/9910505

Orbital Decay in LMC X-4

Alan M. Levine, Saul A. Rappaport, & Goce Zojcheski^{1}

1 Department of Physics and Center for Space Research, Massachusetts Institute of Technology, Cambridge, MA 02139, USA

We report on the results of observations of the binary X-ray pulsar LMC X-4 with the Rossi X-ray Timing Explorer. Our analysis of the Doppler delays of the 13.5-s X-ray pulsations yields the most accurate determination of the LMC X-4 orbital parameters available to date. The epoch of orbital phase zero for the 1.4 day orbit is determined with an uncertainty of ∼ 20 s, and is combined with 5 earlier determinations of the epoch of phase zero to obtain the rate of change in the orbital period: Ṗ_{orb}/P_{orb} = (−9.8 ± 0.7) × 10^{−7} yr^{−1}. This is the first high significance measurement of the rate of change of the orbital period in LMC X-4.

We present data on one of three strong X-ray flares as well as energy-dependent pulse profiles for both non-flaring and flaring time intervals. The pulse profiles during the non-flaring time intervals are typically strikingly different from the flare profiles, but at other times can be similar.

We reconsider the orbital decay of LMC X-4 in the context of tidal evolution. We find that, while the orbital decay is most likely driven by tidal interactions, the asynchronism between the orbit and the rotation of the companion star is probably maintained by the evolutionary expansion of the companion star, just as is thought to be the case for Cen X-3 and SMC X-1. For LMC X-4, we find that the evidence favors the companion star being in a late stage of its life on the main sequence.

Overall, we have found in this work as well as in previous studies that the orbital decay timescales for SMC X-1, Cen X-3, and LMC X-4 are 300,000, 550,000, and 1,000,000 years, respectively. In each case, the Roche lobe will move catastrophically deep into the atmosphere within about 10^{4} years. This short X-ray lifetime may be typical of all massive X-ray binaries in Roche-lobe or near Roche-lobe contact.

Submitted to: The Astrophysical Journal For preprints, contact aml@space.mit.edu

Also available from the URL http://xxx.lanl.gov/abs/astro-ph/9911173

The velocity dispersion of LMC Carbon stars: possible detection of a kinematically distinct population

D.S. Graff, A.P. Gould^{1}, N. Suntzeff, R. Schommer^{2}, E. Hardy^{3}

1 Departments of Astronomy and Physics, The Ohio State University, Columbus, OH 43210, USA 2 Cerro Tololo Inter-American Observatory, Casilla 603, La Serena, Chile 3 National Radio Astronomy Observatory, Casilla 36-D, Santiago, Chile

We analyze the velocity residuals of 551 carbon stars relative to a rotating-disk model of the inner ∼ 70 deg^{2} of the Large Magellanic Cloud (LMC). We find that the great majority of the stars in this sample are best fit as being due to two different populations, a young disk population containing 20% of the stars with a velocity dispersion of 8 km s^{−1}, and an old disk containing the remaining stars with a velocity dispersion of 22 km s^{−1}. The young disk population has a metallicity ∼ 0.25 dex higher than the old disk. However, the data also suggest at the 2σ level that there may be a third kinematically distinct population that is moving towards us at 30 km/sec relative to the LMC, consistent with measurements of 21 cm velocities. If real, this population contains about 7% of the carbon stars in the sample. It could be a feature in the disk of the LMC or it could be tidal debris in the foreground or background. If it is tidal debris, this population could account for some or all of the microlensing events observed towards the LMC.

Submitted to: The Astrophysical Journal

Also available from the URL http://xxx.lanl.gov/abs/astro-ph/9910360

The Star Formation Histories of Two Northern LMC Fields

A. E. Dolphin^{1}

1 National Optical Astronomy Observatories, P.O. Box 26732, Tucson, AZ 85726, USA

Ground-based UBV photometry of two fields in the northern disk of the LMC are presented. A distance modulus of (m−M)_{0} = 18.41±0.04 and an extinction of A_{V} = 0.30±0.05 has been calculated for these fields. The measurable star formation history of the LMC began no more than 12 Gyr ago with a strong star forming episode with [Fe/H] = −1.63 ± 0.10 that accounted for approximately half (by mass) of the LMC’s total star formation in the first 3 Gyr. The data does not give accurate star formation rates during intermediate ages, but there appears to have been a recent increase in the star formation rate in these fields, beginning approximately 2.5 Gyr ago, with the current metallicity in the region being [Fe/H] = −0.38 ± 0.10. The two fields have had very similar star formation rates until 200 Myr ago, at which point one shows a large increase.

Accepted by: Monthly Notices of the Royal Astronomical Society For preprints, contact dolphin@noao.edu

Also available from the URL http://xxx.lanl.gov/abs/astro-ph/9910524

Hubble Space Telescope Photometry of Hodge 301: An “Old” Star Cluster in 30 Doradus

Eva K. Grebel^{1} and You-Hua Chu^{2}

1 University of Washington, Department of Astronomy, Box 351580, Seattle, WA 98195-1580, USA 23 University of Illinois at Urbana-Champaign, Department of Astronomy, 1002 West Green Street, Urbana, IL 61801, USA

We present Hubble Space Telescope Planetary Camera UVI data for the little-studied cluster Hodge 301 3′ northwest of 30 Doradus’ central ionizing cluster R 136. The average reddening of Hodge 301 is found to be ⟨E_{B−V} ⟩ = 0.^{m}28 ± 0.^{m}05 from published infrared and ultraviolet photometry. Using two different sets of evolutionary models, we derive an age of ∼ 20 − 25 Myr for Hodge 301, which makes it roughly 10 times as old as R 136. Hodge 301 is the most prominent representative of the oldest population in the 30 Dor starburst region; a region that has undergone multiple star formation events. This range of ages is an important consideration for the modelling of starburst regions. Hodge 301 shows a widened upper main sequence largely caused by Be stars. We present a list of Be star candidates. The slope of the initial mass function for intermediate-mass main sequence stars ranging from 10 M_{⊙} to 1.3 M_{⊙} is found to be Γ = −1.4 ± 0.1 in good agreement with a Salpeter law. There is no indication for a truncation or change of slope of the IMF within this mass range. In accordance with the age of Hodge 301 no obvious pre-main-sequence stars are seen down to ∼ 1 M_{⊙}. We estimate that up to 41 ± 7 stars with masses > 12M_{⊙} may have turned into supernovae since the formation of the cluster. Multiple supernova explosions are the most likely origin of the extremely violent gas motions and the diffuse X-ray emission observed in the cluster surroundings.

Accepted by: The Astronomical Journal (February 2000 issue)

For preprints, contact grebel@astro.washington.edu
Also available from the URL http://xxx.lanl.gov/abs/astro-ph/9910426

WFPC2 Imaging of Young Clusters in the Magellanic Clouds

Stefan C. Keller, M.S. Bessell & G.S. Da Costa^{1}

Research School of Astronomy and Astrophysics, Institute of Advanced Studies, The Australian National University, Australia

We have carried out Wide Field Planetary Camera 2 F160BW, F555W and F656N imaging of four young populous clusters: NGC 330 in the Small Magellanic Cloud and NGC 1818, NGC 2004 and NGC 2100 in the Large Magellanic Cloud. We report photometric results for these four clusters, including identification using photometric colours of the cluster Be star population. We present theoretical WFPC2 and broad band colours and bolometric corrections for LMC and SMC metallicities. The use of the far-UV F160BW filter enables accurate determination of the effective temperatures for stars in the vicinity of the main sequence turnoff and on the unevolved main sequence.

Accepted by: The Astronomical Journal

For preprints, contact stefan@mso.anu.edu.au
Also available from http://xxx.lanl.gov/abs/astro-ph/9910225

Hubble Space Telescope Observations of the Oldest Star Clusters in the LMC

J.A. Johnson^{1}, M. Bolte^{1}, P.B. Stetson^{2}, J.E. Hesser^{2} and R.S. Somerville^{3}

1 UCO/Lick Observatory, University of California, Santa Cruz, CA 95064, USA 2 Dominion Astrophysical Observatory, Herzberg Institute of Astrophysics, National Research Council of Canada, 5071 West Saanich Road, Victoria, BC V8X 4M6 Canada 3 Racah Institute of Physics, The Hebrew University, Jerusalem, 91904, Israel

We present V, (V − I) color-magnitude diagrams (CMDs) for three old star clusters in the Large Magellanic Cloud (LMC): NGC 1466, NGC 2257 and Hodge 11. Our data extend ∼3 magnitudes below the main-sequence turnoff, allowing us to determine accurate relative ages and the blue straggler frequencies. Based on a differential comparison of the CMDs, any age difference between the three LMC clusters is less than 1.5 Gyr. Comparing their CMDs to those of M92 and M3, the LMC clusters, unless their published metallicities are significantly in error, are the same age as the old Galactic globulars. The similar ages to Galactic globulars are shown to be consistent with hierarchial clustering models of galaxy formation. The blue straggler frequencies are also similar to those of Galactic globular clusters. We derive a true distance modulus to the LMC of (m−M)_{0} = 18.46 ± 0.09 (assuming (m−M)_{0} = 14.61 for M92) using these three LMC clusters.

Accepted by: The Astrophysical Journal

For preprints, contact jennifer@ucolick.org
Also available from the URL http://xxx.lanl.gov/abs/astro-ph/9911187

Final Stages of N-body Star Cluster Encounters

M.R. de Oliveira, E. Bica, H. Dottori

Instituto de Fisica-UFRGS, CP 15051, CEP 91501-970 POA - RS, Brazil

We performed numerical simulations of star cluster encounters with Hernquist’s TREECODE in a CRAY YMP-2E computer. We used different initial conditions (relative positions and velocities, cluster sizes, masses and concentration degrees) with total number of particles per simulation ranging from 4608 to 20480. Long term interaction stages (up to 1 Gyr) when the pair coalesces into a single cluster are compared with isolated LMC clusters. Evidence is found that when seen in a favourable plane these resulting clusters show elliptical shapes due to the disruption of one of the companions. These elliptical shapes are essentially time independent but they do depend on the initial structural parameters of the pair components. We also analyzed the fraction of stars that are ejected to the field by the interaction. We found that this fraction can be almost 50% for the disrupted cluster. These simulations can represent a possible mechanism to explain the ellipticity observed on several star clusters in the Magellanic Clouds.

Accepted by: Monthly Notices of the Royal Astronomical Society For preprints, contact marcio@if.ufrgs.br

Star Formation in a Crossing Time

Bruce G. Elmegreen

IBM T.J. Watson Research Center, P.O. Box 218, Yorktown Heights, NY 10598, USA

Observations suggest that star formation occurs in only one or two crossing times for a range of scales spanning a factor of 1000. These observations include (1) measurements of embedded cluster ages in comparison with the cloud core dynamical times, (2) measurements of the age difference versus separation for clusters in the Large Magellanic Clouds in comparison with the crossing time versus size correlation for molecular clouds, (3) hierarchical structure of embedded young clusters, and (4) the high fraction of dense clouds that contain star formation.

Such a short overall time scale for star formation implies that sources of turbulent energy or internal feedback are not required to explain or extend cloud lifetimes, and that star and protostar interactions cannot be important for the stellar initial mass function. Stars appear in a cloud as if they freeze out of the gas, preserving the turbulent-driven gas structure in their birth locations. The galaxy-wide star formation rate avoids the Zuckerman-Evans catastrophe, which has long been a concern for molecular clouds that evolve this quickly, because the multifractal structure of interstellar gas ensures that only a small fraction of the mass is able to form stars. Star formation on large scales operates slower than on small scales, but in most cases the whole process is over in only a few dynamical times.

Accepted by: The Astrophysical Journal, Vol 530, Feb 10, 2000 For preprints, contact bge@watson.ibm.com

Kepler Rotation Effects on the Binary-Lens Microlensing Events

K. Ioka^{1}, R. Nishi^{1} and Y. Kan-ya^{2}

1 Department of Physics, Kyoto University, Kyoto 606-8502, Japan 2 National Astronomical Observatory, Mitaka, Tokyo, 181-0015, Japan

We investigate the effects of the Kepler rotation of lens binaries on the binary-microlensing events towards the Large Magellanic Cloud (LMC) and the Small Magellanic Cloud (SMC). It is found that the rotation effects cannot always be neglected when the lens binaries are in the LMC disk or the SMC disk, i.e., when they are self-lensing. Therefore we suggest that it will be necessary to consider the rotation effects in the analyses of the coming binary events if the microlensing events towards the halo are self-lensing. As an example, we reexamine the MACHO LMC-9 event, in which the slow transverse velocity of the lens binary suggests a microlensing event in the LMC disk. From a simple analysis, it is shown that the lens binary with total mass ∼ 1M_{⊙} rotates by more than ∼ 60^{∘} during the Einstein radius crossing time. However, the fitting of MACHO LMC-9 with an additional parameter, the rotation period, shows that the rotation effects are small, i.e., the projected rotation angle is only ∼ 5.9^{∘}(M/M_{⊙})^{1/4} during the Einstein radius crossing time. This contradiction can be settled if the physical parameters, such as the mass and the velocity, are different in this event, the binary is nearly edge-on, or the binary is very eccentric, though definite conclusions cannot be drawn from this single event. If the microlensing events towards the halo are due to self-lensing, binary-events for which the rotation effects are important will increase and stronger constraints on the nature of the lenses will be obtained.

Accepted by: Prog. Theor. Phys. Vol.102

For preprints, contact iokakuni@tap.scphys.kyoto-u.ac.jp
Also available from the URL http://xxx.lanl.gov/abs/astro-ph/9910337

Abstracts of Non-Refereed Papers

The Structure of the Circumstellar Envelope of SN 1987A

Arlin P.S. Crotts

Dept. of Astronomy, Columbia University, New York, NY 10027, USA

The volume around the SN 1987A contains a variety of structures, not just the three rings glowing in recombination lines. Many of these are revealed by light echoes, so are mapped in three dimensions by our optical imaging of the SN environs. The rings reside in a bipolar nebula containing them at its waist and crowns, and which is itself contained in a larger, diffuse nebula with a detectable equatorial overdensity. This diffuse nebula terminates in a denser wall which likely marks the inner edge of a bubble blown by the progenitor’s main sequence wind. Along with mapping these structures, we measure spectroscopically the velocity of the gas, revealing, for instance, kinematic ages for the inner and outer rings in close agreement with each other. The presence of these structures, their ages and morphologies must be included in models explaining the evolution of the progenitor star and its mass loss envelope.

To appear in: “Asymmetrical Planetary Nebulae, II: from Origins to Microstructures” ASP Conference Series, J.H. Kastner, N. Soker, & S.A. Rappaport, eds.

For preprints, contact arlin@astro.columbia.edu
Also available from the URL http://xxx.lanl.gov/abs/astro-ph/9910553

Magellanic Cloud Planetary Nebulae: A Fresh Look at the Relations between Nebular and Stellar Evolution

L. Stanghellini^{1,2}, R.A. Shaw^{1}, J.C. Blades ^{1}, and B. Balick^{3}

1Space Telescope Science Institute, 3700 San Martin Drive, Baltimore, MD 21218, USA 2Affiliated to the Astrophysics Division, Space Science Department of ESA 3Astronomy Department, University of Washington, Seattle, USA

Studies of the relationship between planetary nebula morphology and the evolution of the central stars has long suffered from uncertainties in distance determinations, and from the bias of interstellar absorption, that are typical for Galactic PNe. We will be able to eliminate the distance errors and be assured of the sample homogeneity by studying Large Magellanic Cloud (LMC) PNe with images from the Hubble Space Telescope. In this talk we present the first observations in our new sample. The data consist of broad-band images and medium dispersion, slit-less spectra obtained with STIS, and are of excellent quality. Indeed, these data show great promise for subsequent analysis, which will centered on the relationship between nebular morphology and stellar and nebular evolution. While the most intensive analysis of the sample must await the completion of the survey, the data obtained so far show that we will learn a lot along the way.

To appear in: Proceedings of the Conference “Asymmetrical Planetary Nebulae: from Origins to Microstructures” held in Cambridge, MA, August 1999 For preprints, contact lstanghe@stsci.edu

Thesis Abstract

Evolution of large nebulae of the Bar of the Large Magellanic Cloud: Structure of neutral and ionized gas associated.

Patricia Ambrocio-Cruz

IGRAP, Observatoire de Marseille, 2, Place Leverrier, F-13248 Marseille CEDEX 4, France.

Current address: Instituto de Astronomía, UNAM, Apdo. Postal 70-264, México D.F., 04510 México.

We report the study of more than 30 H ii regions in the Large Magellanic Cloud (LMC), in an area where UV photographs show OB associations seen aligned along the stellar bar, and towards the north-west along more than 3.5 kpc. This feature has suggested a spiral arm, and a propagating wave. This site coincides with a position of the LMC in which two H i layers are observed. By the kinematic and dynamic study of nebulae, it is possible to place it in each one of these layers of neutral gas, and to compare to a possible wave triggering the stellar formation in the vicinity of the bar of the LMC. We thus confirm the existence of only two H i layers performing stellar formation, the possible spiral wave being in the low-velocity gas and finishing no farther than N23 to the north.

We have carried out a detailed kinematic study of every one of the nebulae, in order to distinguish the systemic motion from the motions generated by exciting stars. The observations have been obtained with a scanning Fabry-Perot interferometer. For the nebulae of 100-200pc dimension, showing the evolution effects, it was necessary to study the details of their kinematic field and to recall their evolution by comparing the energy input inside the gas with the energy injected by their stellar population. To determine the relations of evolution between the H ii regions observed and their exciting stars, it has been necessary to explore the theoretical models of H ii regions evolving inside a medium with a strong gradient of density.

Knowing the several causes of the detected expansion motions, it is finally possible to discern the structure and the additional effects, induced by other phenomena.

By finding the origin of the kinematic effects, we have deduced that the separation between the two layers decreases from north to south, leading to the encounter point at the place of the N119 nebula. Indeed N119 itself shows characteristics of a rotating system: its rotation and its spiral morphology can be interpreted as resulting from a collision between two molecular clouds coming from every one of the H i layers.

Ph.D. Thesis completed at the Observatoire de Marseille, France, on March, 8^{th} 1999 under the direction of Dr. Annie Laval and Dr. Margarita Rosado.

For copies, contact: patricia@astroscu.unam.mx

Job Opportunity

A postdoctoral position with Dennis Zaritsky at the University of Arizona is open for applications. The postdoc will have access to the Magellanic Clouds Photometric Survey (UBVI photometry of the LMC and SMC). For additional details please see the advertisement posted on the AAS Job Register

or email dzaritsky@as.arizona.edu.